OpenStudy (anonymous):

2x+x=28 complete the square

OpenStudy (anonymous):

is it 2x^2 + x = 28 ????

OpenStudy (anonymous):

yeah sorry

OpenStudy (anonymous):

i no to divide by two and i got x^2+x/2=56

OpenStudy (anonymous):

pls resend the entire question once more as it was given

OpenStudy (anonymous):

2x^2+x=28

OpenStudy (anonymous):

so you divide by two and get x^2+x/2=14

OpenStudy (anonymous):

yes, now you add square of half of the coefficient of x to both sides i. (1/4)^2 x^2 + x/2 + (1/4)^2 = 14 + (1/4)^2

OpenStudy (anonymous):

now left hand side becomes (x+ 1/4)^2 so (x + 1/4)^2 = 14 + 1/16

OpenStudy (anonymous):

225/16

OpenStudy (anonymous):

(x + 1/4)^2 = 225/16 taking squareroot of both sides we get \[x + 1/4 = \pm \sqrt{225/16}\]

OpenStudy (anonymous):

\[x + 14 =\pm \sqrt{225}/\sqrt{16}\]

OpenStudy (anonymous):

where did 14 come from

OpenStudy (anonymous):

prob forgot to put 1/4 small mistake of his.

OpenStudy (anonymous):

soryr made a mistake ....

OpenStudy (anonymous):

its ok just got me confused

OpenStudy (anonymous):

\[x + 1/4 = \pm \sqrt{225}/\sqrt{16}\]

OpenStudy (anonymous):

\[x + 1/4 = \pm15/4\]

OpenStudy (anonymous):

\[x = -1/4 \pm15/4\]

OpenStudy (anonymous):

so you then get x=-1/4+-15/4

OpenStudy (anonymous):

correct and finally x = -1/4 + 15/4 = 14/4 = 7/2 and x = -1/4 - 15/4 = -16/4 = -4

OpenStudy (anonymous):

so we solved the equation by completing the square...

OpenStudy (anonymous):

did u get the method??????